A study was conducted to determine if an expectant mother’s cigarette smoking has an effect on the bone mineral content of her otherwise healthy baby A sample of 77 newborns whose mothers smoked have a mean bone mineral content of Xbar smokers = 0.099 gm/cm with a standard deviation of 0.026 gm/cm. A sample of 161 newborns whose mothers did not smoke have a mean bone mineral content of Xbar nonsmokers = 0.091 gm/cm with a standard deviation of 0.025 gm/cm. Assume that the variance of the underlying populations are equal.

Calculate the t value for the difference between the means.

What is the p value (2 tailed) for the above analysis (hint Excel tdist( t, degrees of freedom, number of tails)
Answer #1

Ans:

pooled standard deviation=sqrt(((77-1)*0.026^2+(161-1)*0.025^2)/(77+161-2))

=0.02533

standard error for difference=0.02533*sqrt((1/77)+(1/161))=0.0035

df=77+161-2=236

Test statistic:

t=(0.099-0.091)/0.0035

t=2.28

p-value(2 tailed)=tdist(2.28,236,2)=0.0235